A)find the quadratic function that models data as a function of x, the number of weeks.
weeks- 0,1,2,3,4,5,6
#- 585,582,629,726,873,1070,1317
B) use the model to estimate the number at 8 weeks

Answer by Edwin McCravy(20054)

A)find the quadratic function that models data as a function of x, the number of weeks.
weeks- 0,1,2,3,4,5,6
#- 585,582,629,726,873,1070,1317

On your TI-83 or TI-84,

Press CLEAR
Press STAT
Press 1

 Put in the values 0,1,2, thru 6, in L1
 Put in the values 585,582, thru 1317, in L2

Press STAT 
Press the right arrow key
Press 5
Press ENTER

You will read:

y=ax²+bx+c
a=25
b=-28
c=585

So quadratic equation (model) is

y = f(x) = 25x² - 28x + 585

B) use the model to estimate the number at 8 weeks

y = f(8) = 25(8)² - 28(8) + 585 = 1961

If you aren't allowed to use a calculator, post again,
stating so.

Edwin